s Question Completion Status: For the following data set, 5,9, 12, 15, 19, 21, 22, 22, 24, 26, 27,27, 31, 33, 33, 34, 37,38, 60,70 a) Find the quartiles 91 Q2 Q3 c) Find the interquartile tange (IQR) IQR- d) Find the lower limit LL = e) Find the upper limit ) identify the potential outiers (Separate each outlier with The potential outlier(s) are

media%2Ffc7%2Ffc7abacd-1c77-47d7-9b82-a0

media%2F67e%2F67e6b264-ed19-4563-97c9-8b
Answer #1

a)

1) First Quartile

n+ 1 20+1 5.25th

Q 1 = 5th item +0.25*(5th -6th)item = 19+0.25*(21-19) =19.5

2) Second Quartile =Median = Q2 = mean of 10th and 11th observation = (26+27)/2 = 26.5

3) Third Quartile

20+115.75th

Q3 = 33+ 0.75(34-33) =33.75

b) 5 number summary:

Min = 5

Q1 = 19.5

Q2 = 26.5

Q3 = 33.75

Max = 70

c) IQR = Q3-Q1 = 33.75-19.5 =14.25

d) Lower Limit = Q1 -1.5IQR = 19.5-1.5*14.25 =-1.875

e) Upper Limit = Q3 +1.5IQR = 33.75+1.5*14.25 =55.125

f) The potential outliers are 60,70